1.    Eight kilometers below the surface of the ocean the pressure is 82Mpa. Determine the density of the seawater at this depth if the density at the surface is 1025kg/m3 and the average bulk modulus of elasticity is 2.3GPa.



1
Expert's answer
2021-08-23T04:49:01-0400

Pressure 8km below the surface:

P2="h\rho g"

=8000*9.81*1025=80.44*106N/m2

"\beta=\frac{∆pV}{∆V}"

"\frac{∆p}{\beta}=\frac{∆V}{V}=1-\frac{\rho}{\rho'}"

"\rho'=\frac{\rho}{1-\frac{∆p}{\beta}}"

∆p=82*106-80.44*106=1.56*106

"\rho'=\frac{1025}{1-\frac{1.56*10^{6}}{2.3*10^{9}}}"

"=\frac{1025}{1-6.7826*10^{-4}}"

"=\frac{1025}{0.9993}"

"=1025.72kg\/m^{3}"